OpenStudy (anonymous):

PLEASE ASSIST ME! i have to sleep already and i need to finish my math project: 6/3x - 18 = -13+2...And i have another problem that may be simpler if thats too complex

OpenStudy (anonymous):

6/3x - 18 = -13+2 first simplify the right side

OpenStudy (anonymous):

15

OpenStudy (anonymous):

-13 + 2 is not 15 ...

OpenStudy (anonymous):

i mean -15

OpenStudy (anonymous):

still not correct....

OpenStudy (anonymous):

if you're having trouble adding positive and negative numbers i suggest you use a number line

OpenStudy (anonymous):

erm what im almost positive

OpenStudy (anonymous):

OHH Sorry

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

6/3x - 18 = -11

OpenStudy (anonymous):

good, now, im assuming the goal is to solve x this means we need to isolate the x term so its alone on one side any ideas what you would do next?

OpenStudy (anonymous):

cancel out -18 by adding it to -11?

OpenStudy (anonymous):

right, you want to cancel out -18 but what would you need to add to BOTH sides in order to cancel out the -18

OpenStudy (anonymous):

6/3x = 7??

OpenStudy (anonymous):

ok yes, you would add 18 to both sides now we have 6/3x = 7 what would you do next?

OpenStudy (nurali):

6/3x = 7 6=7*3x 6=21x x=6/21

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

er, nurali has pointed out an interesting note its the problem (6/3) x =7 or 6/(3x) =7

OpenStudy (anonymous):

hmm im confused sorry

OpenStudy (anonymous):

does the problem look like \[\frac{6}{3} x\] or \[\frac{6}{3x}\]

OpenStudy (anonymous):

the first one

OpenStudy (nurali):

6x/3=7 6x=21 x=21/6

OpenStudy (anonymous):

ok, the next step would be to get rid of the 6/3 from the left side what would you need to divide to both sides to get rid of it

OpenStudy (anonymous):

or nurali will just give the answer thanks alot

OpenStudy (anonymous):

sorry im being slow im just trying to learn step by step i dont understand the last part of how to divide

OpenStudy (anonymous):

6x/3=7 what nurali did was multiply both sides by 3 this gets rid of the 3 in the denominator on the left side this gives us 6x=21 next nurali divided both sides by 6 this gets rid of the 6 on the left sides giving us x=21/6

OpenStudy (anonymous):

ohhh thank you SO MUCH

OpenStudy (anonymous):

yup no problem